DIFFUSE INTERSTELLAR BANDS CORRELATED WITH CARBON MOLECULES C_{2} and C_{3}

J. A. THORBURN, L. M. HOBBS, \underline{T. OKA}, D. E. WELTY and D. G. YORK, Department of Astronomy and Astrophysics, the University of Chicago, Chicago, IL, 60637; B. J. MCCALL, Department of Astronomy, University of California at Berkeley, Berkeley, CA 94720; S. D. FRIEDMAN and P. SONNENTRUCKER, Space Telescope Science Institute, Baltimore, MD 21218; T. P. SNOW, Department of Astronomy, University of Colorado, Boulder CO 80309.

Our survey of diffuse interstellar bands (DIBs) toward 53 stars with color excesses 0.11 ≤ E(B-V) ≤ 1.99 has revealed “the C_{2} DIBs”, a class of about a dozen weak and narrow DIBs whose intensities correlate well with column densities of carbon molecules C_{2} and C_{3} with the correlation coefficient r = 0.85 ∼ 0.50^{a}. They are strongest toward HD 204827, the star whose sightline contains by far the highest C_{2} and C_{3} column densities,^{b} and weak or undetectable toward Herbig’s classic DIB star HD 183143, although the two stars have comparable color excess of 1.11 and 1.27, respectively. The C_{2} DIBs have high correlation among them (r = 0.94 ∼ 0.54, mostly ≥ 0.75) suggesting that their carriers are several molecules with similar chemical properties that exist abundantly in the diffuse interstellar medium where carbon molecules abound. We note among the C_{2} DIBs 4 pairs of doublet lines with very close spacings of 20.9 ∼ 19.1 cm^{−1}. High correlations between the components of a doublet suggest that they are due to the same molecule. The magnitude of the splitting and the relative intensities of the doublets of 4 ∼ 2 suggest that they are spin-orbit split levels of linear molecules.

aJ. A. Thorburn, L. M. Hobbs, B. J. McCall, T. Oka, D. E. Welty, S. D. Friedman, T. P. Snow, P. Sonnentrucker and D. G. York, ApJ \underline{584}, 339 (2003)

bT. Oka, J. A. Thorburn, B. J. McCall, S. D. Friedman, L. M. Hobbs, P. Sonnentrucker, D. E. Welty, and D. G. York, ApJ \underline{582}, 823 (2003)